Let A = {n ∈ Z | n = 5r for some integer r} and B = {m ∈ Z | m = 20s for some integer s}.
a. Is A ⊆ B? Explain.
b. Is B ⊆ A? Explain
1
Expert's answer
2020-11-17T07:37:37-0500

Let "A = \{n\in\mathbb Z\ |\ n = 5r\ \text{for some integer}\ r\}" and "B = \{m\in\mathbb Z\ |\ m = 20s\ \text{for some integer}\ s\}".


a. Since "n=5=5\cdot1" and "1\in\mathbb Z," we conclude that "n=5\in A". On the other hand, "n=5\ne20\cdot s" for each integer "s", and therefore, "n=5\notin B". We conclude that it is not true that "A\subseteq B."


b. If "m\in B" then "m = 20s" for some integer "s". It follows that "m=5(4s), 4s\in\mathbb Z", and consequently, "m\in A". Therefore, "B\subseteq A."